OpenStudy (anonymous):

An equation is shown below: km2 - c = n Which equation is solved for m? m = square root of c plus n over k m = square root of c minus n over k m = the square root of k times the quantity c minus n m = the square root of k times the quantity c plus n

OpenStudy (anonymous):

well, you can easily add c to the both sides and then write : km^2 = n+ m so you have to divide both sides by k: \[m^2 = \frac{ n +m}{ k }\]

OpenStudy (anonymous):

can u say what to do next?

OpenStudy (anonymous):

I dont know where I got the 6 from lol

OpenStudy (anonymous):

yes,@Cosmichaotic is correct,then we should write: \[\sqrt{m^2} = \sqrt{\frac{ m+n }{ k }}\]

OpenStudy (anonymous):